OpenStudy (anonymous):

For the Transformation T, write the T-1. T : (x, y) (5x, 5y ) T -1 (x, y) (?x, ?y)

OpenStudy (anonymous):

I assume you are interested in the inverse.\[T: (x,y)\rightarrow (5x,5y) \implies T^{-1}:(x,y)\rightarrow (\frac{x}{5},\frac{y}{5})\]

OpenStudy (anonymous):

I tried that but It wont accept that answer. It says it's wrong.

OpenStudy (anonymous):

Probably a format issue with the online homework problem. Might need to experiment with different forms, like.... \[(\frac{x}{5},\frac{y}{5})=(\frac{1}{5}x,\frac{1}{5}y)=(.2x,.2y)\]

OpenStudy (anonymous):

Okay I'll try that...thanks

OpenStudy (anonymous):

No sweat. Do math every day.

OpenStudy (anonymous):

Nope, still didnt work.

OpenStudy (anonymous):

Sorry, I don't know what to tell you....

OpenStudy (anonymous):

:/ yeah me neither...i have a friend checking for me. Thanks anyways!